Twenty-five pairs of value of variates X and Y led to the following results: N=25, ∑X=127, ∑Y=100, ∑X^2=760, ∑Y^2=449 and∑XY=500
A subsequent scrutiny showed that two pairs of values were copied down as:
X Y instead of X Y
8 14 8 12
8 6 6 8

Obtain the correct value of the correlation coefficient
1
Expert's answer
2020-03-03T16:48:55-0500

Twenty-five pairs of values of variates X and Y led to the following results: 

"n=25, \sum{X}=127, \sum{Y}=100, \sum{X^2}=760,"

"\sum{Y^2}=449, \sum{XY}=500"

A subsequent scrutiny showed that two pairs of values were copied down as  "\begin{matrix}
 X & Y \\
 8 & 14 \\
 8 & 6
\end{matrix}" while the correct ones are  "\begin{matrix}
 X & Y \\
 8 & 12 \\
 6 & 8
\end{matrix}" . Obtain the correct value of the correlation coefficient. 

New values

"n=25"

"\sum{X}=127-8-8+8+6=125"

"\sum{Y}=100-14-6+12+8=100"

"\sum{X^2}=760-8^2-8^2+8^2+6^2=732"

"\sum{Y^2}=449-14^2-6^2+12^2+8^2=425"

"\sum{XY}=500-8(14)-8(6)+8(12)+6(8)=484"



"r={n\sum{XY}-\sum{X}\sum{Y} \over \sqrt{n\sum{X^2}-(\sum{X})^2}\sqrt{n\sum{Y^2}-(\sum{Y})^2}}"

"r={25\cdot484-125\cdot100 \over \sqrt{25\cdot732-(125)^2}\sqrt{25\cdot425-(100)^2}}\approx-0.309356"